OpenStudy (anonymous):

A student opens a math book to two facing pages. The product of the page numbers is 1640. Find the page numbers. The first page is? The second page is?

OpenStudy (saifoo.khan):

40, 41

OpenStudy (anonymous):

page 1=x page 2=x+1 x(x+1)=1640

OpenStudy (saifoo.khan):

x^2 + x - 1640 solve^

OpenStudy (anonymous):

x^2+x=1640 x^2+x-1640

OpenStudy (anonymous):

poop, saifoo beat me :(

OpenStudy (saifoo.khan):

No Probs man! :)